# Superluminal Wave Activation at Relativistic Magnetized Shocks

### v1

@@ III.2 Dependence of Upstream Scales
where $a_{0}$ is a dimensionless amplitude parameter. We probe the limits of Alfvén wave activation for $\gamma_{\rm u}]^{\rm d}=4$ and an upstream magnetic field of $\mathbf{B}_{\rm u}]^{\rm d}/B_{0}=10$. Figure 2 (left panel) displays a setup that initializes Alfvénic perturbations at kinetic scales in the upstream plasma, using $k_{\rm u}d_{\rm u}]^{\rm u}=0.63$. This implies $\omega_{\rm u}]^{\rm s}/(\bar{\omega}_{\rm pd}/[\gamma_{\rm u}]^{\rm d})\approx 18.37\gg 1$, where we approximate the hot downstream plasma frequency for adiabatic index $\Gamma_{\rm ad}$ and temperature $T$ as
@@
We measure the downstream temperature $T_{\rm d}]^{\rm d}/m_{e}c^{2}=[\left\langle\gamma_{\rm d}\beta_{\rm d}^{2}\right\rangle/3]^{\rm d}=0.96$, where $\left\langle\dots\right\rangle$ averages individual particle velocities. We use the adiabatic index $\Gamma_{\rm ad}=(2+3\,T/m_{e}c^{2})/(1+2\,T/m_{e}c^{2})$ appropriate for two-dimensional Maxwellian distributions, and approximate $T/m_{e}c^{2}$ by the nearest integer. Incoming waves couple to superluminal modes downstream of the shock. For these scales, Equation (15) suggests small amplitudes of residual AWs in the downstream with $\delta B_{A\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 0.03$ for superluminal O-modes propagating into the downstream with a group velocity close to the speed of light. These features are consistently reproduced by the simulation. Two regions form in the downstream: the far downstream region which is reached only by the propagating superluminal O-modes (shaded blue in Figure 2, left panel), and the immediate wake of the shock showing a superposition of superluminal O-modes and non-propagating small amplitude Alfvénic perturbations (shaded red in Figure 2, left panel). In the downstream region, $\mathbf{E}\cdot\mathbf{B}$ is generated at the scale of the propagating O-mode, as would be expected from plasma dispersion (Arons1986; Bransgrove2023).
Figure 2 (right panel) analyzes the same shock and flow properties, but for significantly longer upstream wavelengths $k_{\rm u}d_{\rm u}]^{\rm u}=0.03$. Then, $\omega_{\rm u}]^{\rm s}/(\bar{\omega}_{\rm pd}/[\gamma_{\rm u}]^{\rm d})\approx 0.97\lesssim 1$, and Equation (6) implies no coupling to superluminal O-modes downstream of the shock. The simulation data confirms this prediction. Only non-propagating AWs remain in the immediate downstream, and residual O-mode perturbations with negligible amplitudes and varying wavelengths enter the downstream region. Also, the MHD-scale upstream mode amplitude increases by a factor of approximately two as expected from MHD jump conditions (Appendix A.3).
Figure 3: Wavenumber scaling (top) and frequency matching (bottom) across shocks with $\gamma_{\rm u}]^{\rm d}=4$ (left) and $\gamma_{\rm u}]^{\rm d}=8$ (right). Top panels show downstream wavenumbers as a function of upstream seed wavenumbers in the superposition region for AWs (red circles) and superluminal O-modes (blue circles). The scalings agree closely with Equations (10) and (11). Star markers denote cases with frequencies approximately below the plasma frequency cutoff (Equation 6, indicated by the vertical dashed line) when no propagating downstream modes can be measured. Bottom panels show the dispersion relation of propagating downstream modes. Downstream wavenumbers and corresponding wave frequencies trace the respective (hot) dispersion relation for the SupO-mode with temperatures $T_{\rm d}]^{\rm d}=0.96\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=4$, and $T_{\rm d}]^{\rm d}=2.1\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=8$ (see also Equation 18).
@@ III.2.1 Dispersion and Wave Scales
Figure 3 analyzes simulations for varying upstream wavenumber $k_{\rm u}]^{\rm u}$ and upstream velocity $\gamma_{\rm u}]^{\rm d}$, with the other parameters as in Section III.2. Top panels show Fourier analysis in space for $\gamma_{\rm u}]^{\rm d}=4$ and $\gamma_{\rm u}]^{\rm d}=8$. Bottom panels probe dispersion through temporal FFT at a fixed coordinate in the downstream region. From each simulation, we combine results from four independent FFTs: 1) $E_{y}$ spatially in the supO region (blue circles) 2) $B_{z}$ spatially in the supO/AW region (red circles), 3) $E_{y}$ spatially in the upstream region (red/blue circles) and 4) $E_{y}$ temporally in the supO region to capture the frequency of the downstream propagating modes. We then compare the measured wavenumber to Equations (10) and (11). The predictions capture the downstream wavenumbers well for both non-propagating AWs and propagating modes.
As discussed in Equation (18), estimates of the plasma frequency must consider thermal effects for meaningful normalization of the downstream frequency. We measure downstream temperatures of $T_{\rm d}]^{\rm d}=0.96\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=4$ and $T_{\rm d}]^{\rm d}=2.10\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=8$. This allows us to compare the cold plasma dispersion $T=0$ (dashed blue lines) to the realistic plasma dispersions (dashed green lines) in the bottom panels (b). The measured propagating wave frequencies closely fit the expectation from Equation (6) with corrections for thermal effects. Ultimately, the results from the parameter scan summarized in Figure 3 confirm the frequency and wavenumber matching relationships derived in Section II. We will explore other parameter space dimensions like magnetization and wave-plasma-resonances in the future (see Section IV).
Figure 4: Shock and mode conversion dynamics for a (discrete) spectrum of seed waves. Flow properties are as in Figure 2. We initialize a spectrum of equal amplitude monochromatic waves sampled from $k_{\rm u}d_{\rm u}]^{u}\in\left[0.03,6.28\right]$. We vary the number of modes; the left panel uses the minimum and maximum seed wavelength from Figure 2 ($N=2)$, the middle/right have randomly sampled seed wavenumbers ($N=5$ and $N=10$, in log-space). Low frequency seed waves produce non-propagating downstream AWs (red shaded region). Seed waves with scales above the critical frequency (Equation 6) induce propagating superluminal O-modes (blue shaded region). The bottom panel zooms in on the wave fields $E_{y}]^{\rm d}$ (red color) and $B_{z}]^{\rm d}$ (blue shaded regions).
@@ III.3 Activation of a Spectrum of Seed Perturbations
For this section, we initialize a distribution of seed perturbations resembling turbulent fluctuations upstream of the shock. For a given range of wavenumbers between $k_{\rm u,min}]^{\rm d}$ and $k_{\rm u,max}]^{\rm d}$ we generate a spectrum of $N$ monochromatic waves with random wavenumbers $k_{i,\rm u}]^{\rm d}$ (log-normal distribution), amplitudes $A_{i}$ (Gaussian noise distribution) and phases $\phi_{i}$ (uniform distribution). We normalize individual wave amplitudes with a power-law spectral envelope such that their energy spectrum is approximately $e(k)\propto A(k)^{2}\propto k^{-s}$ for a spectral index $s$ and the total energy across all modes is $E_{\rm tot}$. The wave magnetic field then becomes
$[\delta B_{z,\rm u}]^{\rm d}=\sum_{i=1}^{N}A_{i}\times\cos\left([k_{i,\rm u}]^{\rm d}x+\phi_{i}\right),$ (19)
@@
Figure 4 shows shock and mode conversion dynamics for seed waves sampled of equal amplitude ($s=0$) from $k_{\rm u}d_{\rm u}]^{\rm u}\in\left[0.03,6.28\right]$. For illustration purposes, we choose $N=2$, $N=5$ and $N=10$ superimposed monochromatic modes. The case of $N=2$ combines the two upstream scales displayed in Figure 2. Large scale fluctuations with wavelength above the critical scale defined by Equation (8) become non-propagating AWs downstream of the shock. High-frequency components of the seed spectrum are activated to become propagating O-modes, the shock acts like a high-pass filter. The different components can be identified in the blue (supO) and red (supO/AW) shaded regions of Figure 4. Appendix B discusses spectra of seed waves with $s\approx 1$, with the overall energy distribution of downstream modes following the slope of the seed wave spectrum.
@@ IV Discussion and Conclusions
Ultramagnetized pair-dominated plasmas in compact object magnetospheres allow three different waves: subluminal O-modes (Alfvén waves), superluminal O-modes, and X-modes (Arons1986). Each can have different generation mechanisms and propagation behavior. Alfv́en waves can be generated by plasma instabilities (Usov1987; zeng2025), and magnetic field line motion on the magnetar crust (Carrasco2019; Yuan2020; Yuan2021; Yuan2022; Chen2025; Mahlmann2023; Mahlmann2024). Alfvén waves propagate along magnetic field lines and experience Landau damping, confining their dynamics to the magnetically active inner magnetosphere. Linear X-modes generally do not couple to plasma and can freely propagate in the magnetosphere as EM waves. Long-wavelength X-mode waves can steepen on decaying magnetic fields and generate shocks (chen2022; Beloborodov2023; Vanthieghem2025). Superluminal O-modes do not experience Landau damping and can escape the magnestosphere similar to X-modes. However, plasma properties affect their propagation.
Although superluminal O-modes propagate like an EM wave at high frequencies, they have different polarization than X-modes, with different electric field orientation towards the $\bm{k}-\bm{B}$ plane. Polarimetry observations indicating circular polarization in some FRBs (e.g., Uttarkar2024) could, among other possibilities, suggest that both X- and O-modes contribute to the escaping radio signal. This letter discusses a mechanism of O-mode activation from Alfvénic perturbations at a relativistic shock. In this scenario, Alfvén waves escape the inner magnetosphere by propagating into a magnetized outflow. In the relativistic plasma of the spherically expanding wind, their $k_{\parallel}$ vanishes, essentially ‘freezing’ the Alfvén waves into the flow (see discussion by Thompson2022). At least three types of waves can be generated at the shock: propagating high-frequency superluminal O-modes, low-frequency Alfvénic perturbations, and X-modes via synchrotron maser emission. A first-principles plasma treatment including all interacting modes is essential for a realistic description of radio wave generation at relativistic shocks.
@@ IV.1 Process Efficiency and Frequency Range
to seed propagating downstream modes. Strong magnetized shocks allow for a larger range of upstream fluctuations to be activated. We approximate the resulting downstream scale of superluminal waves via Equation (7):
@@
In this mechanism, interactions of non-propagating AWs with relativistic magnetized shocks drive the activation of superluminal O-modes. Linear coupling generates propagating waves with amplitudes that match the incoming seed waves at scales well above the cutoff frequency in Equation (8). In the analyzed parameter range, we do not find additional coherent generation, resonant enhancement, or suppression of superluminal O-modes. Defining the efficiency of radio wave generation by this process requires the context of global models. It depends on the fraction of available energy that can be transported to the shock location at fluctuation scales that will couple to propagating O-modes in the downstream. The bandwidth of generated superluminal modes is limited at low frequencies by the plasma frequency cutoff in Equation (21). At high frequencies, the signal band extends to the scale of the highest frequency seed fluctuations.
@@ IV.2 Signals in the Observer Frame
Figure 5: Schematic visualization of an observer frame’s view of superluminal O-mode generation via Alfvén wave activation at magnetized shocks. Forward shock (panel a) and reverse shock (panel b) move through a propagating flow, like an expanding magnetar wind, as discussed in Appendix A.2. Depending on the wave propagation direction in the frame of the contact discontinuity (c), amplitude and frequency experience different boosts $\mathcal{C}$, see also Section IV.
@@
Appendix A.2 outlines a simplistic shock structure expected from collisions of homogeneous and highly magnetized ejecta shells (Beloborodov2020; Thompson2022). In such a scenario, the most favorable enhancement of frequencies and amplitudes corresponds to waves generated at the reverse shock (Case A, see also Figure 5). For downstream waves with $\omega_{\rm u}^{2}]^{\rm s}\gg\omega_{\rm pd}^{2}$, wave activation approximately enhances the frequency compared to the upstream perturbation by
@@ IV.3 Astrophysical Implications: Fast Radio Bursts
The luminosity of superluminal O-modes generated by ‘shock-activation’ depends on the available energy in upstream fluctuations above the critical scale given by Equation (20). Thompson2022 discusses the applicability of this model in magnetar magnetospheres. Seed fluctuations generated by activity at the stellar surface likely have spatial scales of the order of a fraction of the neutron star radius $R_{\ast}$, $\lambda_{\rm seed}\approx 0.1R_{\ast}\approx 10^{5}$cm. The plasma skin-depth $d_{0\ast}$ in the immediate magnetar vicinity can be estimated for electron densities $n_{e}=\mathcal{M}n_{\rm GJ}$, where $n_{\rm GJ}=\Omega B_{\ast}/(2\pi ec)$ is the Goldreich-Julian (GJ) density for a magnetosphere rotating with angular frequency $\Omega$ and surface magnetic field $B_{\ast}$. Magnetospheric activity like twists and crustal failures determine the multiplicity $\mathcal{M}$. We find typical values of
@@
Thus, we expect initial seed perturbations with $d_{0\ast}k_{\ast}=2\pi d_{0\ast}/\lambda_{\ast}\lesssim 10^{-7}$. This is likely an upper limit as we expect charge densities above $\mathcal{M}n_{\rm GJ}$ during magnetospheric activities. Modeling the product $k\,d_{0}$ through the magnetosphere and wind zone is not straightforward. The skin-depth $d_{0}$ depends on plasma properties like pair-loading, bulk flow velocity, and temperature, varying substantially between different stages of magnetar activity and magnetospheric models. For example, if expanding in a quasi-thermal pair-fireball, the rest-frame skin-depth scales with radius, roughly $d_{0}\propto r/R_{\ast}$ (cf. Thompson2022). Thus, to exceed the cutoff in Equation (20), the seed waves must cascade over several decades in $k$, reaching $k\,d_{0}\!\approx\!1$. For a cascading spectrum (Appendix B), the energy available at relevant FRB-inducing scales may be limited. For an energy scaling as $e(k)\propto k^{-s}$ for $k>k_{\ast}$, the fraction of available energy above the scale $k_{\rm seed}$ is
@@
For instance, $k_{\rm seed}=10^{3}k_{\ast}$ and $s=5/3$ result in only $\lesssim 1\%$ of the wave energy available for generating propagating waves at the shock with $k>k_{\ast}$.
Observed radio wave properties, like the escaping energy and overall frequency range, could inform magnetosphere models for compact objects like magnetars. Combining seed wave properties with efficiency estimates of transporting fluctuations to the shock location can inform plasma models of these environments. The low-frequency limit on emerging waves (Equation 21) can constrain the plasma magnetization. For the reverse shock scenario (Case A), Equations (20) and (25) estimate a cutoff of the observed frequency,
@@
FRBs typically cover narrowband frequency intervals of a few hundred MHz, in a space between 120MHz and several GHz (PastorMarazuela2021). This range allows for significant variations in the plasma properties contributing to the right-hand side of Equation (28). This complex parameter space requires careful analysis and modeling. The potential effects of baryon loading, particularly in the context of fireball models (bransgrove2025), should also be evaluated.
@@
Activation of Alfvénic perturbations to propagating superluminal modes can apply to other relativistic shocks in magnetically active environments. For magnetars, the plasma properties of relativistic outflows have a large parameter space (e.g., Margalit2018; Beloborodov2020; Thompson2022). Within the light cylinder, large amplitude fast magnetosonic waves can shock when their electric field becomes comparable to the magnetic field (chen2022; Beloborodov2023; Mahlmann2024; Bernardi2025; Vanthieghem2025). Magnetar winds can also form shocks with companion stars in binary systems, possibly associated with periodically repeating radio transients (e.g., Lyutikov2020; Barkov2022; Wang2022; Barkov2024; Wei2024). The described wave activation mechanism has to be evaluated for typical plasma and shock strength parameters of these systems.
@@
Theory for cold pair-plasmas (Section II) and 1D-PIC simulations (Section III) demonstrate the validity of the wave generation mechanism discussed by Thompson2022. Future studies have to evaluate several aspects of this mechanisms, like efficiency of wave activation for varying upstream flow properties. Generalizing to 2D/3D will give crucial insights. Fluctuations at the shock could result in oblique superluminal modes ($k_{\perp}\neq 0$) downstream of the shock. Turbulence fluctuations on scales of the propagating downstream modes could play an important observational role (Thompson2022). Density blobs or magnetic structures could scatter activated superluminal modes and thereby change the properties measured in the observer frame. The coupling between shock-generated (superluminal) plasma waves and downstream turbulence requires numerical and theoretical investigation.

### v2

@@ III.2 Dependence of Upstream Scales
where $a_{0}$ is a dimensionless amplitude parameter and we use $a_{0}=10^{-2}$ in this work. We probe the limits of AW activation for $\gamma_{\rm u}]^{\rm d}=4$ and an upstream magnetic field of $\mathbf{B}_{\rm u}]^{\rm d}/B_{0}=10$. Figure 2 (left panel) displays a setup that initializes Alfvénic perturbations at kinetic scales in the upstream plasma, using $k_{\rm u}d_{\rm u}]^{\rm u}=0.63$. This implies $\omega_{\rm u}]^{\rm s}/(\bar{\omega}_{\rm pd}/[\gamma_{\rm u}]^{\rm d})\approx 18.37\gg 1$, where we approximate the hot downstream plasma frequency for adiabatic index $\Gamma_{\rm ad}$ and temperature $T$ as
@@
We measure the downstream temperature $T_{\rm d}]^{\rm d}/m_{e}c^{2}=[\left\langle\gamma_{\rm d}\beta_{\rm d}^{2}\right\rangle/3]^{\rm d}=0.96$, where $\left\langle\dots\right\rangle$ averages individual particle velocities. We use the adiabatic index $\Gamma_{\rm ad}=(2+3\,T/m_{e}c^{2})/(1+2\,T/m_{e}c^{2})$ appropriate for 2D Maxwellian distributions, and approximate $T/m_{e}c^{2}$ by the nearest integer. Incoming waves couple to superluminal modes downstream of the shock. For these scales, Equation (15) suggests small amplitudes of residual AWs in the downstream with $\delta B_{A\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 0.03$ for superluminal O-modes propagating downstream with a group velocity close to the speed of light. These features are consistently reproduced by the simulation. Two regions form downstream: the far downstream region which is reached only by the propagating superluminal O-modes (shaded blue in Figure 2, left panel), and the immediate wake of the shock showing a superposition of superluminal O-modes and nonpropagating small-amplitude Alfvénic perturbations (shaded red in Figure 2, left panel). In the downstream region, $\mathbf{E}\cdot\mathbf{B}$ is generated at the scale of the propagating O-mode, as would be expected from plasma dispersion (Arons and Barnard, 1986; Bransgrove et al., 2023).
Figure 2 (right panel) analyses the same shock and flow properties, but for significantly longer upstream wavelengths $k_{\rm u}d_{\rm u}]^{\rm u}=0.03$. Then, $\omega_{\rm u}]^{\rm s}/(\bar{\omega}_{\rm pd}/[\gamma_{\rm u}]^{\rm d})\approx 0.97\lesssim 1$, and Equation (6) implies no coupling to superluminal O-modes downstream of the shock. The simulation data confirm this prediction. Only nonpropagating AWs remain in the immediate downstream, and residual O-mode perturbations with negligible amplitudes and varying wavelengths enter the downstream region. Also, the MHD-scale upstream mode amplitude increases by a factor of approximately two as expected from MHD jump conditions (Appendix A.3).
@@ III.2.1 Dispersion and Wave Scales
Figure 3 analyzes simulations for varying upstream wavenumber $k_{\rm u}]^{\rm u}$ and upstream velocity $\gamma_{\rm u}]^{\rm d}$, with the other parameters as in Section III.2. Top panels show Fourier analysis in space for $\gamma_{\rm u}]^{\rm d}=4$ and $\gamma_{\rm u}]^{\rm d}=8$. Bottom panels probe dispersion through temporal FFT at a fixed coordinate in the downstream region. From each simulation, we combine results from four independent FFTs: 1) $E_{y}$ spatially in the supO region (blue circles) 2) $B_{z}$ spatially in the supO/AW region (red circles), 3) $E_{y}$ spatially in the upstream region (red/blue circles) and 4) $E_{y}$ temporally in the supO region to capture the frequency of the downstream propagating modes. We then compare the measured wavenumber to Equations (10) and (11). The predictions capture the downstream wavenumbers well for both nonpropagating AWs and propagating modes.
As discussed in Equation (18), estimates of the plasma frequency must consider thermal effects for meaningful normalization of the downstream frequency. We measure downstream temperatures of $T_{\rm d}]^{\rm d}=0.96\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=4$ and $T_{\rm d}]^{\rm d}=2.10\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=8$. This allows us to compare the cold plasma dispersion $T=0$ (dashed blue lines) to the realistic plasma dispersions (dashed green lines) in the bottom panels (b). The measured propagating wave frequencies closely fit the expectation from Equation (6) with corrections for thermal effects. Estimating the plasma‑frequency cutoff (Equation 6) using the hot‑plasma frequency shifts the activation limit toward higher wavenumbers. In Figure 3, we shade in gray the region between the cutoff frequencies derived for hot and cold plasmas.
The results from the parameter scan in Figure 3 confirm the frequency and wavenumber matching relationships derived in Section II. We will explore other parameter space dimensions like magnetization and wave-plasma resonances in the future (see Section IV).
Figure 4: Shock and mode conversion dynamics for a (discrete) spectrum of seed waves. Flow properties are as in Figure 2. We initialize a spectrum of equal amplitude monochromatic waves sampled from $k_{\rm u}d_{\rm u}]^{u}\in\left[0.03,6.28\right]$. We vary the number of modes; the left panel uses the minimum and maximum seed wavelength from Figure 2 ($N=2)$, the middle/right have randomly sampled seed wavenumbers ($N=5$ and $N=10$, in log space). Low-frequency seed waves produce nonpropagating downstream AWs (red shaded region). Seed waves with scales above the critical frequency (Equation 6) induce propagating superluminal O-modes (blue shaded region). The bottom panel zooms in on the wave fields $E_{y}]^{\rm d}$ (red color) and $B_{z}]^{\rm d}$ (blue shaded regions).
@@ III.3 Activation of a Spectrum of Seed Perturbations
For this section, we initialize a distribution of seed perturbations resembling turbulent fluctuations upstream of the shock. For a given range of wavenumbers between $k_{\rm u,min}]^{\rm d}$ and $k_{\rm u,max}]^{\rm d}$ we generate a spectrum of $N$ monochromatic waves with random wavenumbers $k_{i,\rm u}]^{\rm d}$ (lognormal distribution), amplitudes $A_{i}$ (Gaussian noise distribution), and phases $\phi_{i}$ (uniform distribution). We normalize individual wave amplitudes with a power-law spectral envelope such that their energy spectrum is approximately $e(k)\propto A(k)^{2}\propto k^{-s}$ for a spectral index $s$ and the total energy across all modes is $E_{\rm tot}$. The wave magnetic field then becomes
$[\delta B_{z,\rm u}]^{\rm d}=\sum_{i=1}^{N}A_{i}\times\cos\left([k_{i,\rm u}]^{\rm d}x+\phi_{i}\right).$ (19)
@@
Figure 4 shows shock and mode-conversion dynamics for seed waves sampled of equal amplitude ($s=0$) from $k_{\rm u}d_{\rm u}]^{\rm u}\in\left[0.03,6.28\right]$. For illustration purposes, we choose $N=2$, $N=5$ and $N=10$ superimposed monochromatic modes. The case of $N=2$ combines the two upstream scales displayed in Figure 2. Large-scale fluctuations with wavelength above the critical scale defined by Equation (8) become nonpropagating AWs downstream of the shock. High-frequency components of the seed spectrum are activated to become propagating O-modes, the shock acts like a high-pass filter. The different components can be identified in the blue (supO) and red (supO/AW) shaded regions of Figure 4. Appendix B discusses the spectra of seed waves with $s\approx 1$, with the overall energy distribution of downstream modes following the slope of the seed wave spectrum.
@@ IV Discussion and Conclusions
Ultramagnetized pair-dominated plasmas in compact object magnetospheres allow three different waves: subluminal O-modes (AWs), superluminal O-modes, and X-modes (Arons and Barnard, 1986). Each can have different generation mechanisms and propagation behavior. AWs can be generated by plasma instabilities (Usov, 1987; Zeng et al., 2025), and magnetic field line motion on the magnetar crust (Carrasco et al., 2019; Yuan et al., 2020, 2021, 2022; Chen et al., 2025; Mahlmann et al., 2023, 2024). AWs propagate along magnetic field lines and experience Landau damping, confining their dynamics to the magnetically active inner magnetosphere. Linear X-modes generally do not couple to plasma and can freely propagate in the magnetosphere as EM waves. Long-wavelength X-mode waves can steepen on decaying magnetic fields and generate shocks (Chen et al., 2022; Beloborodov, 2023; Vanthieghem and Levinson, 2025a). Superluminal O-modes do not experience Landau damping and can escape the magnestosphere similar to X-modes. However, plasma properties affect their propagation.
Although superluminal O-modes propagate like an EM wave at high frequencies, they have different polarization than X-modes, with different electric field orientation toward the $\bm{k}-\bm{B}$ plane. Polarimetric observations indicating circular polarization in some FRBs (e.g., Uttarkar et al., 2024) could, among other possibilities, suggest that both X- and O-modes contribute to the escaping radio signal. This letter discusses a mechanism of O-mode activation from Alfvénic perturbations at a relativistic shock. In this scenario, AWs escape the inner magnetosphere by propagating into a magnetized outflow. In the relativistic plasma of the spherically expanding wind, their $k_{\parallel}$ vanishes, essentially ‘freezing’ the AWs into the flow (see discussion by Thompson, 2022). At least three types of waves can be generated at the shock: propagating high-frequency superluminal O-modes, low-frequency Alfvénic perturbations, and X-modes via synchrotron maser emission. A first-principles plasma treatment including all interacting modes is essential for a realistic description of radio wave generation at relativistic shocks.
@@ IV.1 Process Efficiency and Frequency Range
Throughout this discussion, the wavenumber cutoff is defined using the cold plasma frequency. We intentionally adopt the cold plasma approximation to obtain conservative lower limits on the cutoff, since the electron temperature in many of the targeted astrophysical environments is poorly constrained. Hot-plasma corrections generally shift the cutoff to smaller scales (higher wavenumbers). In the simulations presented in Section III, where the temperature is explicitly known, we verified this expectation by benchmarking against the corresponding hot-plasma frequency.
@@
to seed downstream-propagating modes. Strong magnetized shocks allow for a larger range of upstream fluctuations to be activated. We approximate the resulting downstream scale of superluminal waves via Equation (7):
@@
In this mechanism, interactions of nonpropagating AWs with relativistic magnetized shocks drive the activation of superluminal O-modes. Linear coupling generates propagating waves with amplitudes that match the incoming seed waves at scales well above the cutoff frequency in Equation (8). In the analyzed parameter range, we do not find additional coherent generation, resonant enhancement, or suppression of superluminal O-modes. Defining the efficiency of radio wave generation by this process requires the context of global models. It depends on the fraction of available energy that can be transported to the shock location at fluctuation scales that will couple to propagating O-modes in the downstream. The bandwidth of generated superluminal modes is limited at low frequencies by the plasma-frequency cutoff in Equation (21). At high frequencies, the signal band extends to the scale of the highest-frequency seed fluctuations.
@@ IV.2 Signals in the Observer Frame
Figure 5: Schematic visualization of an observer frame’s view of superluminal O-mode generation via Alfvén wave activation at magnetized shocks. A forward shock (panel a) and reverse shock (panel b) move through a propagating flow, like an expanding magnetar wind, as discussed in Appendix A.2. Depending on the wave propagation direction in the frame of the contact discontinuity, amplitude and frequency experience different boosts $\mathcal{C}$, see also Section IV.
@@
Appendix A.2 outlines a simplistic shock structure expected from collisions of homogeneous and highly magnetized ejecta shells (Beloborodov, 2020; Thompson, 2022). In such a scenario, the most favorable enhancement of frequencies and amplitudes corresponds to waves generated at the reverse shock (Case A, see also Figure 5). For downstream waves with $\omega_{\rm u}^{2}]^{\rm s}\gg\omega_{\rm pd}^{2}$, wave activation approximately enhances the frequency compared to the upstream perturbation by
@@ IV.3 Astrophysical Implications: Fast Radio Bursts
The luminosity of superluminal O-modes generated by ‘shock-activation’ depends on the available energy in upstream fluctuations above the critical scale given by Equation (20). The fraction of turbulent energy available at $kd_{0}\sim 1$ may be very small for realistic spectra and scale separations and is the main source of uncertainty when applying our results to FRB sources. Thompson (2022) discusses the applicability of the shock-activation model in magnetar magnetospheres. Seed fluctuations generated by activity at the stellar surface likely have spatial scales of the order of a fraction of the neutron star radius $R_{\ast}$, $\lambda_{\rm seed}\approx 0.1R_{\ast}\approx 10^{5}$cm. The plasma skin depth $d_{0\ast}$ in the immediate magnetar vicinity can be estimated for electron densities $n_{e}=\mathcal{M}n_{\rm GJ}$, where $n_{\rm GJ}=\Omega B_{\ast}/(2\pi ec)$ is the Goldreich-Julian (GJ) density for a magnetosphere rotating with angular frequency $\Omega$ and surface magnetic field $B_{\ast}$. Magnetospheric activity like twists and crustal failures determine the multiplicity $\mathcal{M}$. We find typical values of
@@
Thus, we expect initial seed perturbations with $d_{0\ast}k_{\ast}=2\pi d_{0\ast}/\lambda_{\ast}\lesssim 10^{-7}$. This is likely an upper limit as we expect charge densities above $\mathcal{M}n_{\rm GJ}$ during magnetospheric activities. Modeling the product $k\,d_{0}$ through the magnetosphere and wind zone is not straightforward. The skin depth $d_{0}$ depends on plasma properties like pair loading, bulk flow velocity, and temperature, varying substantially between different stages of magnetar activity and magnetospheric models. For example, if expanding in a quasi-thermal pair fireball, the rest-frame skin depth scales with radius, roughly $d_{0}\propto r/R_{\ast}$ (cf. Thompson, 2022). Thus, to exceed the cutoff in Equation (20), the seed waves must cascade over several decades in $k$, reaching $k\,d_{0}\!\approx\!1$. For a cascading spectrum (Appendix B), the energy available at relevant FRB-inducing scales may be limited. For an energy scaling as $e(k)\propto k^{-s}$ for $k>k_{\ast}$, the fraction of available energy above the scale $k_{\rm seed}$ is
@@
For instance, $k_{\rm seed}=10^{3}k_{\ast}$ and $s=5/3$ result in only $\lesssim 1\%$ of the wave energy available for generating propagating waves at the shock with $k>k_{\ast}$. The small fraction of turbulent energy at $kd_{0}\sim 1$ makes quantitative predictions for FRBs uncertain.
Observed radio wave properties, like the escaping energy and overall frequency range, could inform magnetosphere models for compact objects like magnetars. Combining seed wave properties with efficiency estimates of transporting fluctuations to the shock location can inform plasma models of these environments. The low-frequency limit on emerging waves (Equation 21) can constrain the plasma magnetization. For the reverse shock scenario (Case A), Equations (20) and (25) estimate a cutoff of the observed frequency:
@@
FRBs typically cover narrowband frequency intervals of a few hundred MHz, in a space between 120MHz and several GHz (Pastor-Marazuela et al., 2021). This range allows for significant variations in the plasma properties contributing to the right-hand side of Equation (28). The complex parameter space requires careful analysis and modeling. The potential effects of baryon loading, particularly in the context of fireball models (Bransgrove et al., 2025), should also be evaluated.
@@
Activation of Alfvénic perturbations to propagating superluminal modes can apply to other relativistic shocks in magnetically active environments. For magnetars, the plasma properties of relativistic outflows have a large parameter space (e.g., Margalit and Metzger, 2018; Beloborodov, 2020; Thompson, 2022). Within the light cylinder, large-amplitude fast magnetosonic waves can shock when their electric field becomes comparable to the magnetic field (Chen et al., 2022; Beloborodov, 2023; Mahlmann et al., 2024; Bernardi et al., 2025; Vanthieghem and Levinson, 2025a). Magnetar winds can also form shocks with companion stars in binary systems, possibly associated with periodically repeating radio transients (e.g., Lyutikov et al., 2020; Barkov and Popov, 2022; Wang et al., 2022; Barkov et al., 2024; Wei et al., 2024). The described wave activation mechanism has to be evaluated for typical plasma and shock strength parameters of these systems.
@@
Theory for cold pair plasmas (Section II) and 1D PIC simulations of linear, low-amplitude, seed perturbations (Section III) demonstrate the validity of the wave generation mechanism discussed by Thompson (2022). Future studies have to evaluate several aspects of this mechanism, like efficiency of wave activation for varying upstream flow properties. The activation criterion discussed in this letter applies to Alfvénically polarized, nonpropagating magnetic shear fluctuations that are advected into the shock. Generalizing to 2D/3D will give crucial insights. If the upstream fluctuations were propagating AWs with finite $k_{\parallel}$, the associated electric fields could affect the mode matching across the shock and add resonant interactions with the shock structure. The activation condition derived here should therefore be viewed as the nonpropagating limit and a conservative threshold for more general upstream conditions. Fluctuations at the shock could result in oblique superluminal modes ($k_{\parallel}\neq 0$) downstream of the shock. Turbulence fluctuations on scales of the downstream-propagating modes could play an important observational role (Thompson, 2022). Density blobs or magnetic structures could scatter activated superluminal modes and thereby change the properties measured in the observer frame. The coupling between shock-generated (superluminal) plasma waves and downstream turbulence requires numerical and theoretical investigation.